OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

OpenStudy (anonymous):

sin3x=sin(2x+x)=sin2x*cosx+sinx*cos2x can you continue??

OpenStudy (anonymous):

ahhhhh is it sin x cos2x - sin3x + cos3x

OpenStudy (anonymous):

no it is'nt. do you know that sin(a+b)=sina*cosb+sinb*cosa

OpenStudy (anonymous):

no. 2 sin x cos2x + sin3x

OpenStudy (anonymous):

answer is \[2sinx*\cos ^{2}x+sinx-2\sin ^{3}x \] or; \[3sinx*\cos ^{2}x-\sin ^{3}x\] or; \[4sinx*\cos ^{2}x-sinx\] ok??